OpenStudy (anonymous):

Sven drove the 119 km from Bournemouth to Bristol at an average speed of 68 km/h.Work out the time, in hours and minutes, his journey took. The answer is not important, rather the workings out.

OpenStudy (chaise):

t = d/s t = 119/68 time = 119/68 2.4791666666666666666 ~2.48 48x0.60 = 28.8 2 hours and ~29 mins,

OpenStudy (anonymous):

119KMx 1H/68KM= 1.75 H 1.75 H X 60min/ 1H= 105min

OpenStudy (anonymous):

the answer is 1 hour 45 minutes

OpenStudy (anonymous):

i got that from my textbook, so are those workings correct?

OpenStudy (chaise):

You're right - for some reason instead of dividing by 68 like I wrote down, I divided by 48, which explains the dodgy answer. Very sorry -.-

OpenStudy (anonymous):

Yes, its a conversion question. They may ask for the question in a specific format.

OpenStudy (anonymous):

@ chaise what is 2.48??

OpenStudy (chaise):

Scroll up - Unintentionally divided by 48 rather than 68, which gave a dodgy answer.

OpenStudy (anonymous):

would you mind writing it out again for me please?

OpenStudy (anonymous):

Brilliant at watsgood

OpenStudy (anonymous):

@ chrissy if you look at the answer I gave you see the conversion factor and how they cancel out. You either get the ans. in Hours or minutes you can't use the units simultaneously. You can combine as the final answer... is what you see in the textbook.

OpenStudy (anonymous):

bloody brilliant, thank you as well @chaise